What is the range of values of signed integers using a 4 byte representation?
Answer #1
Number of bits for 4 bytes = 4*8 = 32

Minimum value of signed integers using a 4 byte representation = -2^31 = -2147483648
Minimum value of signed integers using a 4 byte representation = (2^31)-1 = 2147483647

So, Range of values of signed integers using a 4 byte representation is from -2147483648 to 2147483647